OpenStudy (anonymous):

sqrt(x+6) = sqrt(x – 10)-3

OpenStudy (anonymous):

\(\sqrt{x+6} = \sqrt{x-7} -3\)

OpenStudy (anonymous):

You need to start by squaring both sides of the equation. \[(\sqrt{x+6})^2 = (\sqrt{x-7} - 3)^2\]

OpenStudy (anonymous):

\[x+6 = (\sqrt{x-7} - 3)(\sqrt{x-7}-3)\]

OpenStudy (anonymous):

Expand and then isolate the remaining square-root on one side of the equation. Square both sides again. Then, isolate x.

OpenStudy (anonymous):

can you please do one more step

OpenStudy (anonymous):

\[x+6 = (x-7) -3\times\sqrt{x-7} - 3\times\sqrt{x-7}+9\] \[x+6 = x-7 - 6\times\sqrt{x-7} + 9\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

However....no solutions exist.

OpenStudy (anonymous):

Did you type the problem correctly, or is that a possible answer (multiple choice)? Else, if this is hand-written homework, I would just write that there are no solutions.

OpenStudy (anonymous):

does this one has a solution

OpenStudy (anonymous):

sqrt(x+6) = sqrt(x – 10)-3

OpenStudy (anonymous):

thank once again

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

Thanks. :)